# 3D Visual Illusion Depth Estimation

### v2

@@ Abstract
3D visual illusion is a perceptual phenomenon where a two-dimensional plane is manipulated to simulate three-dimensional spatial relationships, making a flat artwork or object look three-dimensional in the human visual system. In this paper, we reveal that the machine visual system is also seriously fooled by 3D visual illusion, including monocular and binocular depth estimation. In order to explore and analyze the impact of 3D visual illusion on depth estimation, we collect a large dataset containing almost 3k scenes and 200k images to train and evaluate SOTA monocular and binocular depth estimation methods. We also propose a robust depth estimation framework that uses common sense from a vision-language model to adaptively select reliable depth from binocular disparity and monocular depth. Experiments show that SOTA monocular, binocular, and multi-view depth estimation approaches are all fooled by various 3D visual illusions, while our method achieves SOTA performance. https://github.com/YaoChengTang/3D-Visual-Illusion-Depth-Estimation
@@ 1 Introduction
Depth estimation is a task of recovering 3D geometry of a given image or image sequence. It is a long-standing and unresolved vision problem with a lot of researches, like monocular depth estimation, stereo matching, and multi-view reconstruction. These works have achieved impressive perfromance in common scenes, almost can rival that of human perception. Despite the common scenes, there are also many challenging scenes that make a flat artwork or object look three-dimensional in the visual system, like 3D video advertisements and transparent glass, as shown in Figure 1. These challenging scenes contain 3D visual illusion, which will result in wrong depth perception and seriously affect the downstream applications, casuing safety-critical risks in AR/VR and robotics.
In this paper, we present a 3D-Visual-Illusion dataset to explore the impact of 3D visual illusion on depth estimation. The dataset contains 5 kinds of illusions: inpainting illusion (e.g., inpainting on a wall/floor), picture illusion (e.g., picture printed/drawn on a paper), replay illusion (e.g., video replayed on different screens), holography illusion, and mirror illusion (e.g., virtual image on reflective objects). The dataset is collected in almost 3k scenarios with 200k images, across from indoors to outdoors, from small objects to large street views. We build the dataset using virtual data and real-world data. The virtual data is generated from web-source videos and generative models using two kind of data generation pipelines. The real-world data is collected using a stereo camera and a solid-state LiDAR depth camera. The evalution of SOTA depth estimation models on our dataset show that the monocular priors learned from large-scale data is easily fooled by illusion textures, like pictures and replayed videos, while the matching results fail at special materials, like transparent glass and mirrors.
Figure 1: The visualization of 3D visual illusion.
@@
Inspired by the strong generalization ability of vision-language models (VLMs) across diverse tasks, we introduce a VLM-driven monocular-stereo fusion model to address the challenges posed by 3D visual illusions. The proposed model leverages the commonsense knowledge from VLM to assess the reliability of monocular and binocualr depth in different regions, enabling effective depth fusion. The model consists of a dual-branch prediction network and a VLM-based fusion network. The dual-branch prediction network takes a rectified image pair as input and uses two branches to predict monocular depth and binocular disparity simultaneously. The VLM-based fusion network first use a pre-trained large vision model to extract features from monocular depth, binocular disparity, and RGB image. The features are further transformed into a shared embedding space with language prompt using a large language model. The embedding features are then used to generate confidence map via flow matching. The confidence map is used to guide the alignment of mnocular depth from affine-invariant space to metric space. The aligned monocular depth is then fused with the binocular disparity to generate the final depth map. The experiments on our dataset and Booster dataset show that the proposed method achieves SOTA performance on different 3D visual illusions.
@@ 2.1 Stere Matching
Stereo matching is a pixel-wise labeling task that relies on dense correspondence between a pair of images. The SOTA methods are either GRU-based iterative methods or Transformer-based methods. The former methods predict the disparity update and iteratively approximate the GT value in a GRU framework [26, 23, 16, 45, 5, 10]. They have achieved great performance in both benchmark and zero-shot generalization testing. Teh latter methods use Transformer to learn matching and predict the disaprity map [24, 12, 40, 41, 38, 22, 36]. The Transformer-based methods achieve superior performance by learning from a large-scale data. In this paper, we collect a comprehensive, large-scale dataset to thoroughly investigate and evaluate the impact of 3D visual illusions on matching methods. Our findings reveal that stereo matching methods are highly susceptible to various illusions, while our method leverage common sense from a large vision-language model to help rectify these illusions.
@@ 2.2 Monocular Depth Estimation
Monocular depth estimation is a pixel-wise regression task based on single image. Recent deep learning methods leverage diffusion models [17, 8, 47] or Transformers [31, 2, 44, 42] to extract depth-related features, in both supervised and self-supervised settings [9, 11, 19]. Despite their impressive generalization performance across diverse scenes, these methods fundamentally rely on monocular cues, which are susceptible to 3D visual illusions, much like the human visual system. In this work, we introduce a large-scale benchmark to evaluate state-of-the-art monocular depth models under such illusions. Our results show that existing methods are consistently misled by these challenging patterns. To address this, we propose leveraging matching-based depth cues as a complementary information to enhance monocular depth estimation.
@@ 2.3 Large Vision-Language Model
The large vision-language model (VLM) injects common sense from billions of textual data to support vision understanding and generation [46]. It presents great power in various task, like visual question answering, image generation, and navigation. The methods of these tasks mainly adapt a pre-trained VLM to specific datasets to preserve the generalization ability, while promoting the understanding of specific tasks. To further facilitate the training of VLM in downstream tasks, a lot of methods explore the different finetuning strategy, like Prompt [4], Adapter [14], LoRA [15], and LST [33]. In this paper, we propose to use VLM to predict the confidence of the disparity map to recover the metric version of monocular depth. The common sense from large VLM is beneficial for the confidence estimation in various complex scenes.
@@
We construct a 3D-Visual-Illusion dataset to explore the challenging of 3D visual illusion in the depth estimation tasks. The 3D-Visual-Illusion dataset is constructed in almost 3k scenarios with 178912 frames used for training and 617 frames used for testing. The dataset contains images with various resolutions with the maximum resolution of $1080\times 1920$ across from indoors to outdoors, from small objects to large street inpainting. It contains 5 kinds of illusions: inpainting illusion (e.g., inpainting on a wall/floor), picture illusion (e.g., picture printed/drawn on a paper), replay illusion (e.g., video replayed on different screen), holography illusion, and mirror illusion (e.g., virtual image on reflective objects). The 3D-Visual-Illusion dataset is generated/captured in the virtual world and the real world respectively. We will introduce the details of construction for the virtual data subset and the real data subset in the following contents.
@@ 3.1 Virtual Data
We collect a large amount of video data from websites and text-to-video generative models. We take the videos as left image sequence and generate disaprity maps and right images.
Video Collection  We adopt two distinct data collection strategies for the first four types of illusions and for mirror illusions. For inpainting, picture, replay, and holography illusions, we crawl 5,226 web videos (over 52M frames) using keyword-based search. We then use a large vision-language model Qwen2-VL-72B [1, 37] to automaticly delete bad frames, reducing the dataset to 4,519 videos (1.4M frames). Then, we manually remove blurry or occluded frames, resulting in 1,384 high-quality videos (236K frames). For mirror illusions, due to the rarity of mirror-related keywords, mirror illusion videos are hard to find from the web. Instead, we employ generative models (Sora [28], Kling [21], and HunyuanVideo [20]) to generate videos. Prompts are initially generated with ChatGPT and iteratively refined. After manual filtering, we collect 234 high-quality mirror illusion videos containing 2,382 frames.
Depth Generation  After collecting videos from both web and generative models, we generate depth for each frame using the pipelines shown in Figure 2 and 3. We use different data generation pipelines for the two kinds of data due to the following reasons: (1) Web-sourced data mainly use fixed cameras, offering limited viewpoints and making accurate scene reconstruction challenging. (2) Videos from generative models are used for mirror illusions, requiring the geometry of the mirror world for right image generation. However, monocular depth estimation may predict the geometry of either the mirror surface or the reflected scene, making it difficult to control.
For web videos, we use the pre-trained DepthAnything V2 [43] to predict inverse depth, interpreted as disparity under unknown camera parameters. However, the generated disparity in 3D illusion regions is completely wrong. To address this, we introduce a neighboring support region as a reference to correct the disparity of illusion regions by assuming they are located on the same plane. Segmentation masks for both illusion and support regions are obtained using SAM2 [32]. Due to the limitations of automatic mode in detecting illusions, we manually annotate all frames using SAM2’s click mode, filtering out redundant or imperceptible cases. This yields 176,530 high-quality annotated frames.
@@
After obtaining the mask of illusion and support regions, a plane can be fitted using the points within a support region. In the standard 3D camera coordinate system $(X,Y,Z)$, the general form of a plane with parameters $(\alpha,\beta,\gamma,\delta)$ is $\alpha\cdot x+\beta\cdot y+\gamma\cdot z+\delta=0$. Given the relationship between image coordinate $(u,v)$, disparity $d$, and $(X,Y,Z)$:
@@
the planar structures in 3D space $(X,Y,Z)$ remain planar in the disparity space $(u,v,d)$ with plane parameters $(\alpha,\beta,\delta,\gamma)$: $\alpha\cdot u+\beta\cdot v+\delta\cdot d+\gamma=0$. This is crucial for our generation, as it allows us to fit a 3D plane using the disparity values under unknown camera baseline and intrinsics, without converting disparity into depth. Given a set of $N$ points from the support region, $\{(u_{i},v_{i},d_{i})\}_{i=1}^{N}$, the goal of plane fitting becomes a least squares fitting problem about parameters $(\alpha,\beta,\delta,\gamma)$. To mitigate the impact of noise during fitting, we adopt RANSAC [7] for robust plane fitting (For more details, please refer to the supplemental materials). The fitted plane is then used to rectify disparity values within illusion regions. After obtaining the rectified disparity map, we further denoise the disparity to accquire smooth boundaries between support regions and their surroundings.
For videos from generative model, we reconstruct the entire scene using InstantSplat [6], which first derives geometry from DUSt3R [38] and generates novel views using GS [18]. We extract disparity information from the DUSt3R geometry and further refine it through a series of post-processing steps, similar to aforementioned semantic segmentation, RANSAC plane fitting, and disaprity denoising.
Right Image Generation  The right-view images for generative-model videos are directly rendered using Gaussian Splatting (GS). For web-sourced videos, right views are synthesized by warping the left images using monocular disparity. Due to scale ambiguity, we estimate an optimal scale factor $s$ via binary search, terminating when most warped pixels fall within the valid image width:
@@ 3.2 Real-world Data
In addition to virtual data, we collect real-world data comprising 72 scenes and 617 frames. The setup includes a stereo camera (ZED Mini) and a LiDAR-based depth camera (Realsense L515). Details of the setup are provided in the supplemental materials. To ensure accurate alignment, the two cameras are rigidly mounted, and their relative pose is calibrated using a checkerboard. Based on the calibration, the L515 depth map is warped to the ZED left camera to construct GT depth.
@@
$z_{Z}^{*}=\min_{(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}},z_{Z}^{{}^{\prime}})\in P_{m}}\{Z_{Z}^{{}^{\prime}}\mid$ $(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}})=(u_{Z},v_{Z})\},$
Although upsampling greatly densifying point cloud, projecting from a lower-resolution to a higher-resolution space may still introduce small holes. To address this, we apply connected component analysis to identify missing regions and fill them using image inpainting [35]. The inpainted depth values are further smoothed to ensure seamless transitions with surrounding areas.
@@
$\mathcal{Z}_{Z}(u_{Z},v_{Z})$ =$=\begin{cases}0&\text{if }\mathcal{Z}_{L}(u_{Z\rightarrow L},v_{Z\rightarrow L})==0\ or\\ &\quad\left|z_{Z\rightarrow L}-\mathcal{Z}_{L}(u_{Z\rightarrow L},v_{Z\rightarrow L})\right|>\tau,\\ \mathcal{Z}_{z}(u_{z},v_{z})&\text{otherwise}.\end{cases}$ (6)
Here, $\tau$ is a manually defined threshold. To further refine depth quality, median filtering is applied to suppress noise and remove outlier points from the point cloud. Finally, the depth map $\mathcal{Z}_{Z}$ is converted into disparity map $D$:
@@ 4 VLM-Driven Monocular-Stereo Fusion Model
We draw inspiration from the strong generalization ability of vision-language models (VLMs) across diverse tasks to address the 3D-visual-illusion challenge. We propose a VLM-driven monocular-stereo fusion model that leverages commonsense knowledge to assess the reliability of monocular and stereo depth in different regions, enabling effective depth fusion. As illustrated in Figure 4, our model contains two components: a dual-branch prediction network and a VLM-based fusion network.

### v4

@@ Abstract
3D visual illusion is a perceptual phenomenon where a two-dimensional plane is manipulated to simulate three-dimensional spatial relationships, making a flat artwork or object look three-dimensional in the human visual system. In this paper, we reveal that the machine visual system is also seriously fooled by 3D visual illusions, including monocular and binocular depth estimation. In order to explore and analyze the impact of 3D visual illusion on depth estimation, we collect a large dataset containing almost 3k scenes and 200k images to train and evaluate SOTA monocular and binocular depth estimation methods. We also propose a 3D visual illusion depth estimation framework that uses common sense from the vision language model to adaptively fuse depth from binocular disparity and monocular depth. Experiments show that SOTA monocular, binocular, and multi-view depth estimation approaches are all fooled by various 3D visual illusions, while our method achieves SOTA performance.
[Uncaptioned image]    [Uncaptioned image]    [Uncaptioned image]
@@ 1 Introduction
Depth estimation aims to recover the 3D geometry of a scene from a single image or an image sequence. It is a long-standing and challenging vision problem, with extensive research in monocular depth estimation [46, 47, 3, 17], stereo matching [27, 42, 5, 13], and multi-view reconstruction [41, 39]. These works have achieved impressive performance in typical, well-structured scenes, approaching human-level perception. However, beyond typical scenes, there exist many 3D visual illusion scenes that make a flat artwork or object look three-dimensional, as illustrated in Figure 1. These 3D visual illusions mislead the depth perception and seriously affect the downstream applications, causing safety-critical risks in AR/VR and robotics.
Figure 1: The visualization of 3D visual illusions.
In this paper, we present a 3D-Visual-Illusion dataset to investigate the impact of 3D visual illusions on depth estimation. The dataset includes five types of illusions: inpainting illusion (e.g., inpainting on walls or floors), picture illusion (e.g., image printed/drawn on a paper), replay illusion (e.g., videos replayed on different screens), holography illusion, and mirror illusion (e.g., specular and transparent surfaces). It comprises nearly 3,000 scenes and 200,000 images, covering various environments from small objects to large scenes and from indoor to outdoor settings. We construct the dataset from both virtual and real-world data. Virtual data is generated using two separate pipelines: one based on web-sourced videos and the other on generative models. Real-world data is captured using a stereo camera and a solid-state LiDAR depth sensor.
The evaluation results on the 3D-Visual-Illusion dataset reveal distinct failure modes for different SOTA depth estimation models. Monocular methods, which rely on the mapping from texture cues to 3D geometry, are easily misled by illusion patterns such as printed images or screen content. In contrast, stereo methods depend on pixel correspondences and fail on transparent or reflective surfaces like glass and mirrors, where conflicting signals distort the matching process. Notably, stereo and monocular methods exhibit complementary strengths, often succeeding where the other fails. Stereo methods succeed on texture-rich illusions, while monocular models can recover mirror geometry through learned priors. Yet, each alone is insufficient to handle the full spectrum of 3D visual illusions, and this complementarity motivates us to seek a unified framework that leverages the strengths of both.
@@
Inspired by the strong generalization capability of vision-language models (VLMs) on mirror illusions (see the supplementary materials for details), we propose a VLM-driven monocular–stereo fusion framework to fuse stereo and monocular priors. The model leverages commonsense knowledge from VLMs to assess the reliability of monocular and binocular depth across different regions, enabling more effective depth fusion. Our model consists of two components: a dual-branch prediction network and a VLM-based fusion network. The dual-branch network takes a rectified image pair as input and simultaneously predicts monocular depth and binocular disparity. The VLM-based fusion network employs a pre-trained vision model to extract features from the left RGB image. These features are mapped into a shared embedding space using a large language model conditioned on a language prompt. The embedding features are then used to generate a confidence map via flow matching. The confidence map is used to align the affine-invariant monocular depth to metric scale, which is then fused with the binocular disparity to produce the final depth map. Experiments on our dataset and the Booster dataset demonstrate that our method achieves SOTA performance under a wide range of 3D visual illusions.
@@ 2.1 Stere Matching
Stereo matching is a pixel-wise labeling task that relies on dense correspondence between a pair of images. The SOTA methods are either GRU-based iterative methods or Transformer-based methods. The former methods predict the disparity update and iteratively approximate the GT value in a GRU framework [27, 24, 16, 48, 5, 10]. They have achieved great performance in both benchmark and zero-shot generalization testing. The latter methods use a Transformer to learn matching and predict the disparity map [25, 12, 43, 44, 41, 23, 39]. The Transformer-based methods achieve superior performance by learning from large-scale data. In this paper, we collect a comprehensive, large-scale dataset to thoroughly investigate and evaluate the impact of 3D visual illusions on matching methods. We reveal that stereo matching methods are highly susceptible to various illusions. Our method leverages common sense from VLM to detect mirror illusions and help rectify these illusions.
@@ 2.2 Monocular Depth Estimation
Monocular depth estimation is a pixel-wise regression task based on a single image. Recent deep learning methods leverage diffusion models [17, 8, 50] or Transformers [33, 2, 47, 45] to extract depth-related features, in both supervised and self-supervised settings [9, 11, 19]. Despite their impressive generalization performance across diverse scenes, these methods fundamentally rely on monocular cues, which are susceptible to 3D visual illusions, much like the human visual system. In this work, we introduce a large-scale benchmark to evaluate state-of-the-art monocular depth models under such illusions. Our results show that existing methods are consistently misled by these challenging patterns. To address this, we propose leveraging matching-based depth cues as complementary information to enhance monocular depth estimation.
@@ 2.3 Large Vision-Language Model
The large vision-language model (VLM) injects common sense from billions of textual data to support vision understanding and generation [49]. It presents great power in various tasks, like visual question answering, image generation, and navigation. The methods of these tasks mainly adapt a pre-trained VLM to specific datasets to preserve the generalization ability, while promoting the understanding of specific tasks. To further facilitate the training of VLM in downstream tasks, a lot of methods explore different finetuning strategies, like Prompt [4], Adapter [14], LoRA [15], and LST [35]. In this paper, inspired by the strong detection ability of VLM on mirror illusions, we use VLM to predict the confidence of the disparity map to recover the metric version of monocular depth. The common sense from large VLM is beneficial for the confidence estimation in various complex scenes.
@@
We construct the 3D-Visual-Illusion dataset to investigate the challenges posed by 3D visual illusions in depth estimation. The dataset comprises nearly 3,000 scenarios, with over 200,000 frames for training and 617 frames for testing. It includes images of various resolutions, up to a maximum of $1080\times 1920$, and spans a wide range of scenes, from indoor environments and small objects to large-scale street views. The dataset covers five types of illusions: inpainting illusion (e.g., inpainting on a wall/floor), picture illusion (e.g., picture printed/drawn on a paper), replay illusion (e.g., video replayed on a different screen), holography illusion, and mirror illusion (e.g., specular or transparent surfaces). Data is collected from both virtual and real-world sources. Details of the construction process for the virtual and real subsets are provided in the following sections.
@@ 3.1 Virtual Data
We collect a large amount of video data from websites and text-to-video generative models. We take the videos as left image sequences and generate disparity maps and right images.
Video Collection  We adopt two distinct data collection strategies for the first four types of illusions and for mirror illusions. For inpainting, picture, replay, and holography illusions, we crawl 5,226 web videos (over 52M frames) using keyword-based search. We then apply a vision-language model, Qwen2-VL-72B [1, 40], to automatically filter out irrelevant frames, reducing the dataset to 4,519 videos (1.4M frames). Further manual filtering removes blurry or occluded frames, resulting in 1,384 high-quality videos with 236K frames. Mirror illusions are difficult to collect from the web due to the rarity of mirror-related keywords and high-quality videos. To address this, we generate videos using SOTA generative models, including Sora [30], Kling [21], and HunyuanVideo [20]. Prompts are initially created with ChatGPT and manually refined. Videos violating physical plausibility are discarded. In total, we collect 234 high-quality mirror illusion videos comprising 2,382 frames.
@@
Depth Generation  After collecting videos from both web sources and generative models, we generate depth for each frame using the pipelines illustrated in Figure 2 and 3. Different pipelines are used for the two data sources for the following reasons: (1) Web-sourced videos typically involve fixed cameras, providing limited viewpoints and making accurate scene reconstruction difficult. (2) Generative videos are used primarily for mirror illusions, which require modeling the geometry of the reflected (mirror) world to generate right-view images. However, monocular depth estimation may ambiguously predict either the mirror surface or the reflected scene, leading to inconsistent results.
For web videos, we use the pre-trained DepthAnything V2 [46] to predict inverse depth, which is treated as disparity under unknown camera parameters. However, in regions affected by 3D visual illusions, the predicted disparity is often severely inaccurate. To correct this, we introduce a neighboring support region, assuming it lies on the same plane as the illusion region, and use it as a reference for disparity correction. Segmentation masks for both illusion and support regions are obtained using SAM2 [34]. Since the automatic mode struggles to detect illusions accurately, we manually annotate all frames using SAM2’s click mode, removing redundant or imperceptible cases.
After obtaining the mask of illusion and support regions, we fit a plane using the points within a support region. In the standard 3D camera coordinate system $(X,Y,Z)$, the general form of a plane with parameters $(\alpha,\beta,\gamma,\delta)$ is $\alpha\cdot x+\beta\cdot y+\gamma\cdot z+\delta=0$. Given the relationship between image coordinate $(u,v)$, disparity $d$, and $(X,Y,Z)$:
@@
the planar structures in 3D space $(X,Y,Z)$ remain planar in the disparity space $(u,v,d)$ with plane parameters $(\alpha,\beta,\delta,\gamma)$: $\alpha\cdot u+\beta\cdot v+\delta\cdot d+\gamma=0$. This property is crucial for our generation, as it allows plane fitting directly in disparity space under unknown camera intrinsics and baseline, avoiding the need to convert disparity into depth. Given a set of $N$ points from the support region, $\{(u_{i},v_{i},d_{i})\}_{i=1}^{N}$, the goal of plane fitting becomes a least squares fitting problem over parameters $(\alpha,\beta,\delta,\gamma)$. To mitigate the impact of noise during fitting, we adopt RANSAC [7] for robust plane estimation (see supplemental materials for details). The fitted plane is then used to rectify disparity values within the illusion regions. After obtaining the rectified disparity map, we apply an additional denoising step to ensure smooth transitions along the boundaries between support regions and their surroundings.
For videos from generative models, we reconstruct the entire scene using InstantSplat [6], which first estimates geometry via DUSt3R [41] and synthesizes novel views using Gaussian Splatting (GS) [18]. We extract the disparity map from the geometry derived from DUSt3R and refine it through a series of post-processing steps, including semantic segmentation, RANSAC-based plane fitting, and disparity denoising.
Right Image Generation  The right-view images for generative-model videos are directly rendered using Gaussian Splatting (GS). For web-sourced videos, right views are generated by warping the left images using monocular disparity. Due to scale ambiguity, we estimate an optimal scale factor $s$ via binary search, terminating when most warped pixels fall within the valid image width:
@@ 3.2 Real-world Data
In addition to virtual data, we collect real-world data comprising 72 scenes and 617 frames. The setup includes a stereo camera (ZED Mini) and a LiDAR-based depth sensor (Realsense L515), with details provided in the supplemental materials. To ensure accurate alignment, the two sensors are rigidly mounted, and their relative pose is calibrated using a checkerboard. The L515 depth map is then warped to the ZED left camera frame based on the calibration to construct the ground-truth depth.
@@
$z_{Z}^{*}=\min_{(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}},z_{Z}^{{}^{\prime}})\in P_{m}}\{Z_{Z}^{{}^{\prime}}\mid$ $(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}})=(u_{Z},v_{Z})\}.$
Although upsampling greatly densifies the point cloud, projecting from a lower-resolution to a higher-resolution space may still introduce small holes. To address this, we apply connected component analysis to identify missing regions and fill them using image inpainting [38]. The inpainted depth values are further smoothed to ensure seamless transitions with surrounding areas.
@@
[…]
Here, $\epsilon$ is a manually defined threshold. To further refine depth quality, median filtering is applied to suppress noise and remove outlier points from the point cloud. Finally, the depth map $\mathcal{Z}_{Z}$ is converted into disparity map $D$:
@@ 4 VLM-Driven Monocular-Stereo Fusion Model
As illustrated in Figure 4, our model first uses a dual-branch prediction network to predict the binocular disparity and the monocular depth. Then, a VLM-based fusion network is used to produce the final disparity map by fusing the binocular disparity and the monocular depth.